# When Precision Meets Position: BFloat16 Breaks Down RoPE in Long-Context Training

### v1

@@ 2.2 BFloat16 Disrupts RoPE’s Relative Properties, Especially with Long Context
$D(\mathbf{X},\Delta_{1},\Delta_{2})=\sum_{l,h}\sum_{t_{\text{row}}}\mathbf{n}\odot\left(\sum_{t_{\text{col}}}\left|\text{ATTN}^{l,h}(\mathbf{X},\Delta_{1})-\text{ATTN}^{l,h}(\mathbf{X},\Delta_{2})\right|\right),$ (4)
This metric measures the cumulative difference in attention scores across all layers and heads. The $t_{\text{row}}$ and $t_{\text{col}}$ iterate over the rows and columns of the attention matrix. And the normalization vector $\mathbf{n}=\left[\frac{1}{L},\frac{1}{L-1},\dots,1\right]$ is applied element-wise (denoted by $\odot$) to account for the varying number of elements in lower triangular matrices due to causal masking.
@@ 4.2 Controllable Study with Meaningful Evaluations
However, unlike the official evaluation, which measures overall performance across all 13 tasks, we exclude two tasks (NIAH-Single 3 and NIAH-Multikey 3) for the LLaMA-2-7B model. This exclusion is based on the assumption that long-context training primarily extends a model’s existing abilities from its original context window to longer sequences, and on the observation that LLaMA-2-7B struggled with these two tasks even within its original 4K context window. It is therefore unreasonable to expect the model to acquire new skills in extended contexts that it could not perform within its original window. (Detailed results for both LLaMA-2-7B and LLaMA-2-7B-Chat across all 13 tasks along with a comprehensive discussion are provided in Appendix D). Note that for other advanced models (LLaMA-3-8B, Mistral-7B-v0.3, Qwen-1.5-1.8B) in the cross-architecture evaluation section (Section 5.3), we use all 13 tasks.
@@ Appendix D LLaMA-2-7B RULER Performance
In Table 10, we use the base template for all models. Compared to the results reported in Hsieh et al. (2024), we observed that Llama 2 7B’s performance on the Common Word Extraction (CWE) task is suboptimal. This occurs because when the context length is set to 4,000 (less than 4,096), an example task will be provided (refer to the generate_input_output function in scripts/data/synthetic/common_words_extraction.py in the RULER Hsieh et al. (2024). As shown in Table 11, when we set the context length to $4,096$, we can reproduce the reported result. The inclusion of an example disrupts the Llama 2 7B model’s ability to accurately count words, reflecting the instability of the CWE task for evaluating Llama 2 7B. Therefore, we omit this task from the long context evaluation.
Additionally, based on our hypothesis that continued training of large language models primarily enhances their ability to handle longer contexts rather than introducing new capabilities, we have excluded the NIAH Multikey 3, QA 1, and QA 2 tasks from our evaluation. This decision stems from the observation that even within the pretraining context length, LLaMA 2 7B is unable to effectively solve these tasks.

### v2

@@ 2.2 BFloat16 Disrupts RoPE’s Relative Properties, Especially with Long Context
$D(\mathbf{X},\Delta_{1},\Delta_{2})=\sum_{l,h}\sum_{j=1}^{L}\mathbf{n}\odot\left(\sum_{i=1}^{L}\left|\text{ATTN}^{l,h}_{i,j}(\mathbf{X},\Delta_{1})-\text{ATTN}^{l,h}_{i,j}(\mathbf{X},\Delta_{2})\right|\right),$ (4)
This metric measures the cumulative difference in attention scores across all layers and heads. And the normalization vector $\mathbf{n}=\left[\frac{1}{L},\frac{1}{L-1},\dots,1\right]$ is applied element-wise (denoted by $\odot$) to account for the varying number of elements in lower triangular matrices due to causal masking.
@@ 4.2 Controllable Study with Meaningful Evaluations
However, unlike the official evaluation, which measures overall performance across all 13 tasks, we exclude two tasks (and NIAH-Multikey 3 and Common Word Extraction) for the LLaMA-2-7B model. This exclusion is based on the assumption that long-context training primarily extends a model’s existing abilities from its original context window to longer sequences, and on the observation that LLaMA-2-7B struggled with these two tasks even within its original 4K context window. It is therefore unreasonable to expect the model to acquire new skills in extended contexts that it could not perform within its original window. (Detailed results for both LLaMA-2-7B and LLaMA-2-7B-Chat across all 13 tasks along with a comprehensive discussion are provided in Appendix D). Note that for other advanced models (LLaMA-3-8B, Mistral-7B-v0.3, Qwen-1.5-1.8B) in the cross-architecture evaluation section (Section 5.3), we use all 13 tasks.
@@ Appendix D LLaMA-2-7B RULER Performance
In Table 10, we use the base template for all models. Compared to the results reported in Hsieh et al. (2024), we observe that Llama 2 7B’s performance on the Common Word Extraction (CWE) task is unstable. When the context length is set to 4,000 (less than 4,096), a slightly different context example is provided (refer to the generate_input_output function in scripts/data/synthetic/common_words_extraction.py in RULER (Hsieh et al., 2024)). As shown in Table 11, when the context length is set to $4,096$, we can reproduce the reported results. However, the inclusion of a slightly different example disrupts the Llama 2 7B model’s ability to accurately count words, highlighting the instability of the CWE task for evaluating Llama 2 7B. Consequently, we omit this task from the long-context evaluation.
Furthermore, based on our hypothesis that continued training of large language models primarily improves their ability to handle longer contexts rather than introducing new capabilities, we have excluded the NIAH Multikey 3 and CWE tasks from our evaluation. This decision is supported by the observation that even within the pretraining context length, Llama 2 7B is unable to effectively solve these tasks.